# Expected Points Above Average: A Novel NBA Player Metric Based on Bayesian Hierarchical Modeling

### v1

@@ 6 Conclusions and Discussion
In this research we develop a Bayesian hierarchical model to cluster NBA teams and players based on their shot taking tendencies and shot making abilities. Our approach enables full probabilistic inference of important team- and player-based metrics, including expected points and a novel expected points above average metric. The methodology is highlighted using team and player based comparisons from the 2020-21 season.
@@
We note that our work on expected points and EPAA line up with each of the three goals of sports analytics research as outlined in the introduction: evaluating individual impact, improving winning, and quantifying components of the game. In particular, EPAA is a novel method to evaluate an individual basketball player’s scoring, which enables teams to better identify players who can contribute to their team’s success. Both metrics represent novel methods for quantifying shooting in the sport of basketball.
The NBA game has naturally evolved over time with the most obvious change being a greater emphasis on three-point shooting in recent years. We utilized the clustering methodology described in Section 3.2 to account for changes, as well as commonalities, in shooting characteristics over time. We should mention, again, that the goal of this work was to use the posterior samples to develop new basketball metrics and not on inference with respect to estimated clusters themselves.
@@
In addition, our work provides an additional way to evaluate individual impact via the EPAA metric. The application could improve a team’s overall quality by allowing team management to identify the best offensive player and quantifying components of the game by examining team shooting profiles over time. From a management perspective, this type of analysis would help to identify players who have similar performance metrics yet are paid at markedly different rates.
In the spirit of openWAR (Baumer, Jensen and Matthews, 2015), we stress that the data and results are fully open and reproducible. In particular, we provide a github repository (https://github.com/rtelmore/EPAA) and a Shiny web application for additional exploration and extensions of our work.
@@
Finally, we believe that this methodology and our new metrics, Expected Points and Expected Points Above Average, could easily be extended to other sporting domains. For example, one could imagine applying this methodology and similar concepts to data in the National Football League (NFL) to cluster teams and potentially develop novel player-specific metrics, e.g. passing yards above average for quarterbacks. Furthermore, shooting and/or passing tendencies amongst soccer teams and players in the English Premier League (EPL), Major League Soccer (MLS), or other leagues is a natural extension. Likewise, a similar analysis could shed more insight into playing tendencies in the National Hockey League (NHL).

### v2

@@ 6 Conclusions and Discussion
In this research we develop a Bayesian hierarchical model to cluster NBA teams and players based on their shot taking tendencies and shot making abilities. Our approach enables full probabilistic inference of important team- and player-based metrics, including expected points and a novel expected points above average metric. The methodology is highlighted using team- and player-based comparisons from the 2020-21 season.
@@
The NBA game has naturally evolved over time, with the most obvious change being a greater emphasis on three-point shooting in recent years. We utilized the clustering methodology described in Section 3.2 to account for changes, as well as commonalities, in shooting characteristics over time. Through the use of posterior samples, the contribution of this work was to develop new basketball metrics. Importantly, expected points and EPAA aligns with each of the three goals of sports analytics research as outlined in the introduction: evaluating individual impact, improving winning, and quantifying components of the game. In particular, EPAA is a novel method to evaluate the scoring of an individual basketball player, which enables teams to better identify players who can contribute to their success. Both metrics represent novel methods for quantifying shooting in the sport of basketball.
@@
In addition, the EPAA metric provides a novel way to evaluate individual impact. The application could improve a team’s overall quality by allowing team management to identify the best offensive player and quantifying components of the game by examining team shooting profiles over time. From a management perspective, this type of analysis would help to identify players who have similar performance metrics yet are paid at markedly different rates.
@@
We believe that this methodology and our new metrics, Expected Points and Expected Points Above Average, could easily be extended to other sporting domains. For example, this methodology and similar concepts could be applied to data in the National Football League (NFL) to cluster teams and potentially develop novel player-specific metrics, e.g. passing yards above average for quarterbacks. Furthermore, shooting and/or passing tendencies among soccer teams and players in the English Premier League (EPL), Major League Soccer (MLS), or other leagues are natural extensions. Likewise, a similar analysis could shed more insight into playing tendencies in the National Hockey League (NHL).
Despite these benefits of our proposed model and inference, there are a few limitations to this work. First, placing NBA (or any sport) players on an ”average” team is an impossible task in practice. However, we believe, as do many practitioners, that it offers insight into a new dimension of individual player performance. Basketball is a team game, and metrics such as EPAA ignore potential synergistic effects between teammates as they relate to scoring. Additionally, under different nuisance parameter settings, we observed slight variations in the posterior distributions of expected points per game and EPAA for teams and players, respectively. Importantly, however, this had little impact on conclusions drawn from the results of our model given that the ranking of the teams and players remained fairly consistent across these specifications.
There are several possible directions for extending this work. Specifically, estimating the number of both shot taking and accuracy clusters in a data-driven manner is ripe for exploration. The problem is well known in the literature on finite mixture models, see, for example, McLachlan (2000). Unlike the Bayesian approach offered here in which Dirichlet priors are used, the utilization of sparse finite mixture models to infer the number of clusters is becoming increasingly popular (Frühwirth-Schnatter and Malsiner-Walli, 2019; Malsiner-Walli, Frühwirth-Schnatter and Grün, 2016). A comparison of this approach with our current analytical pipeline would certainly be an interesting study.